Show that the coefficients of the given differential equation are homogeneous and solve the given differential equation: "-ydx+(x+\u221axy)dy=0"


1
Expert's answer
2022-02-21T16:19:16-0500

"-ydx+(x+\sqrt{xy})dy=0\\
(x+\sqrt{xy})dy=ydx\\
y'=\frac{y}{x+\sqrt{xy}} \equiv f(x,y)\\
f(tx,ty)=\frac{ty}{tx+\sqrt{tx\cdot ty}}=\\
=\frac{ty}{t(x+\sqrt{x y})}=\frac{y}{x+\sqrt{xy}}=f(x,y)\\
 y=x\cdot z(x), y'=x\cdot z'+z\\
xz'+z=\frac{xz}{x+\sqrt{x\cdot xz}}\\
xz'+z=\frac{z}{1+\sqrt{z}}\\
xz'=\frac{z}{1+\sqrt{z}}-z\\
xz'=\frac{-z\cdot\sqrt{z}}{1+\sqrt{z}}\\
\frac{1+\sqrt{z}}{z\cdot\sqrt{z}}dz=-\frac{dx}{x}\\
\int(\frac{1}{z^{\frac{3}{2}}}+\frac{1}{z})dz=-\int\frac{dx}{x}\\
-2 z^{-\frac{1}{2}}+\ln|z|=-\ln|x|-\ln|C|\\
2 z^{-\frac{1}{2}}-\ln|z|=\ln|x|+\ln|C|\\
\frac{e^{2z^{-\frac{1}{2}}}}{z}=x\cdot C"

return to y

"\frac{e^{2(\frac{y}{x})^{-\frac{1}{2}}}}{\frac{y}{x}}=x\cdot C\\
e^{2(\frac{y}{x})^{-\frac{1}{2}}}=Cy."